Solve the first order linear inhomogeneous differential equation using the constant variation method

y,- (3y/x)=x


1
Expert's answer
2022-02-15T18:20:36-0500

Corresponding homogeneous differential equation



"y'-\dfrac{3y}{x}=0""\dfrac{dy}{y}=3\dfrac{dx}{x}"

Integrate



"\int \dfrac{dy}{y}=\int 3\dfrac{dx}{x}""\ln (|y|)=3\ln (|x|)+\ln C""y=Cx^3"

The general solution of the homogeneous differential equation is



"y_h=Cx^3"

Use the constant variation method



"y'=C'x^3+3x^2 C"

Substitute



"C'x^3+3x^2 C-\dfrac{3Cx^3}{x}=x""C'=x^{-2}"

Integrate



"C=\int x^{-2}dx""C=-\dfrac{1}{x}+C_1"

Finf the general solution of the non homogeneous differential equation



"y=(-\dfrac{1}{x}+C_1)x^3""y=-x^2+C_1x^3"